# Spectro-polarimetric view of the gamma-ray emitting NLS1 1H0323+342

### v1

@@ 2 Observations and data reduction
We observed 1H0323+342 using SPOL (Schmidt et al., 1992) mounted on 6.5m MMT on 20th December 2019. Observations were carried out using a slit of 1.5 arcsec, and the 964 l/mm grating providing a spectral resolution of 15 Å in the wavelength range of 4100–7200Å to cover both the H and H regions. A rotatable semi-achromatic half-wave plate was employed to modulate the incoming polarization. At the same time, a Wollaston prism, having very high throughput and polarimetric efficiency, in the collimated beam separates the orthogonally polarized spectra onto a thinned and anti-reflection-coated 800 × 1200 SITe CCD. Four separate exposures, capturing 16 orientations of the wave plate, were conducted. This provided us with background-subtracted measures of each normalized linear Stokes parameter, q and u. In total, 1 hour of on-source observation was made to obtain high signal-to-noise ratio (S/N) data.
@@
The polarization angles are adjusted to ensure they fall within the range of 0 to 180 degrees, using the formula and conditions provided by Bagnulo et al. (2009).
@@ 3.2.1 Effect of Inter-stellar polarization
Capetti et al. (2021) corrected the interstellar polarization for their sources based on the Planck polarization images. Recently, this method has been described in detail for IRAS 04416+1215 (AV=1.187) by Śniegowska et al. (2023). We followed the same method and calculated interstellar polarization based on I, Q, and U Planck images at frequency 353 GHz from skyview444https://skyview.gsfc.nasa.gov/current/cgi/query.pl. The images were re-sampled, smoothing by 9 pixels. The details of the calculation are given in Capetti et al. (2021) and Śniegowska et al. (2023). Interstellar polarization is found to be 0.22% in the V-band, which is relatively high. We subtract the ISP in the Stokes (Q, U) parameter space to correct the interstellar polarization contribution from the data. The corrected polarization degree and position angle are % with degrees over the whole wavelength range, % and degrees in the continuum, as mentioned in the Table 3. The degree of polarization after correcting for interstellar polarization has been increased twice.
@@ 4 Modeling polarized emission using STOKES
In our investigation of the scattering region representing the inner-edge of torus, which was modeled as a flared disk with defined inner and outer radii and a half-opening angle. We assumed that the scattering is primarily caused by free electrons, with the number density parameter governing its properties. Similar to the approach taken for the BLR, we set the velocity distribution, assuming consistency in the inner and outer radii of the torus. The inner radius is determined using the R-L relation derived from infrared reverberation mapping studies (Kishimoto et al., 2007; Koshida et al., 2014). Following Figure 2 in Savić et al. (2018), the outer radius for the scattering region is chosen such that the BLR half-opening angle when viewed from the edge of the scattering region is 35o. This angle gives the angular extension of the BLR while keeping the observer at the outer edge of the scattering region. We choose a slightly larger value of the half-opening angle compared to Savić et al. (2018, where they assumed a BLR half-opening angle = 25o) which allows us to reproduce well the wings of the emission lines. This value was obtained after testing with a grid of values between 0o - 45o (with a step size of 5o; where 0o refers to no contribution from the BLR and 45o means that the outer edge of the BLR coincides with the outer edge of the scattering region). Values above 45o are discarded as they are nonphysical. Therefore, with the knowledge of the from the scaling relation of Netzer & Laor (1993), and this value of the BLR half-opening angle, we can estimate the outer radius for the scattering region ( = ). By combining information on inner and outer radii with electron number density, we estimate the optical depth of this region. A grid of electron number density solutions is explored to identify the optimal solution aligning with spectral data gathered from observations.
@@ 5 Discussions
We conducted spectro-polarimetric observations of the gamma-ray-emitting NLS1, 1H0323+342, renowned for its strong Fe II emission, with an R4570 of 1.14 and accretion at 37% of the Eddington limit. The continuum polarization is modest, measuring at only 0.12%, aligning with the observed low-polarization trend in other high-accreting NLS1 sources. Goodrich (1989) performed spectro-polarimetry of 18 NLS1s, majority of which show a very low level of polarization %. Robinson et al. (2011) utilized a two-component scattering model, determining that the inclination of a line-emitting accretion disk is not a principal parameter, suggesting the distinctive properties of these objects have a physical rather than geometrical origin. Indeed, recent studies on high-accreting NLS1 show higher metallicities in these objects (e.g., Panda et al., 2019, 2020; Śniegowska et al., 2021; Panda, 2021; Garnica et al., 2022).
@@ 5.4 Inferences about the inclination and geometry
The inter-percentile width mentioned in Table 4 can be used to calculate the inclination. The implies and the disk thickness must be high enough to explain the low-inclination angle inferred from the STOKES modeling. Assuming the inclination angle degrees implies a = 0.8 or disk half-opening angle of 40 degrees (). A relatively higher inclination angle of 24 degrees will correspond to a H/R = 0.7 and, consequently, a disk half-opening angle of 35 degrees. Indeed, recent dynamical modeling of reverberation mapping data, e.g., Williams et al. (2018), supports the argument that BLR is a thick disk where many objects show a half-opening angle () larger than 40 degrees (here =90°corresponds to a sphere). The BLR can have flared disk-like geometry where increases with R as (see discussion in Collin et al., 2006). Such flaring ‘BOWL’-shaped BLR models can explain various observational signatures, including the mechanism responsible for producing Lorentzian profiles in Balmer lines in low-inclination systems dominated by turbulence (see Goad et al., 2012).
@@ Acknowledgement
We thank the referee for her/his valuable comments and suggestions, which helped to improve the quality of the manuscript. The authors thank Paul Smith for his valuable support in the observations and SPOL data reduction and Alessandro Capetti for his help in the estimation of inter-stellar polarization. JJ and SR acknowledge the partial support of SERB-DST, New Delhi, through SRG grant no. SRG/2021/001334. SP acknowledges the financial support of the Conselho Nacional de Desenvolvimento Científico e Tecnológico (CNPq) Fellowships 300936/2023-0 and 301628/2024-6. NS acknowledges the financial support provided under the National Post-doctoral Fellowship (NPDF; Sanction Number: PDF/2022/001040) by the Science & Engineering Research Board (SERB), a statutory body of the Department of Science & Technology (DST), Government of India. Observations reported here were obtained at the MMT Observatory, a joint facility of the University of Arizona and the Smithsonian Institution.

### v2

@@ 2 Observations and data reduction
We observed 1H0323+342 using SPOL (Schmidt et al., 1992) mounted on 6.5m MMT on 20th December 2019. Observations were carried out using a slit of 1.5 arcsec, and the 964 l/mm grating, providing a spectral resolution of 15 Å in the wavelength range of 4100–7200Å to cover both the H and H regions. A rotatable semi-achromatic half-wave plate was employed to modulate the incoming polarization. At the same time, a Wollaston prism, having very high throughput and polarimetric efficiency, in the collimated beam separates the orthogonally polarized spectra onto a thinned and anti-reflection-coated 800 × 1200 SITe CCD. Four separate exposures, capturing 16 orientations of the wave plate, were conducted. In each orientation, the ratio of fluxes of the two orthogonally polarized spectra was used to calculate the polarization measurement. This provided us with background-subtracted measures of each normalized linear Stokes parameter, q and u (see Schmidt et al., 1992; Smith et al., 2003). In total, 1 hour of on-source observation was made to obtain high signal-to-noise ratio (S/N) data.
@@
Following Wardle & Kronberg (1974), we attempted debiasing the degree of polarization (see also Eq. 1 of Smith et al., 2009) to estimate the statistical bias inherent in these measurements. However, the correction was negligible due to binning the spectra (see section 3.2). We found that debiasing changes the results of the degree of polarization by only, which is insignificant. Therefore, this polarization debiasing is not considered for further analysis. The polarization angles are adjusted to ensure they fall within the range of 0 to 180 degrees, using the formula and conditions provided by Bagnulo et al. (2009).
@@ 3.2.1 Effect of Inter-stellar polarization
Capetti et al. (2021) corrected the interstellar polarization for their sources based on the Planck polarization images. Recently, this method has been described in detail for IRAS 04416+1215 (AV=1.187) by Śniegowska et al. (2023). We followed the same method and calculated interstellar polarization based on I, Q, and U Planck images at frequency 353 GHz from skyview444https://skyview.gsfc.nasa.gov/current/cgi/query.pl. The images were re-sampled, smoothing by 9 pixels. The submillimeter polarization vector (, ) is converted to the visible using and where the ratio of submillimeter to V-band polarization = (stat.) 0.3 (syst.), fairly constant across different lines of sight in the diffuse interstellar medium (see Planck Collaboration et al., 2015), and is the absorption optical depth. Then, the empirical law from Serkowski et al. (1975) is applied to convert the V-band ISP polarization to the wavelength of the H line using 5500 as the reference point at which the ISP reaches its maximum (). The calculated ratio between and , is 0.79. The errors in the I, Q, and U Planck maps are not propagated when calculating errors in ISP polarization. Further details of the calculation are given in Capetti et al. (2021) and Śniegowska et al. (2023). Interstellar polarization is found to be 0.22% in the V-band, which is relatively high. We subtract the ISP in the Stokes (Q, U) parameter space to correct the interstellar polarization contribution from the data. The corrected polarization degree and position angle are % with degrees over the whole wavelength range, % and degrees in the continuum, as mentioned in the Table 3. The degree of polarization after correcting for interstellar polarization has been increased twice.
@@ 4 Modeling polarized emission using STOKES
In our investigation of the scattering region representing the inner edge of the torus, we modeled it as a flared disk with defined inner and outer radii and a half-opening angle. We assumed that the scattering is primarily caused by free electrons, with the number density parameter governing its properties. Similar to the approach taken for the BLR, we set the velocity distribution, assuming consistency in the inner and outer radii of the torus. The inner radius is determined using the R-L relation derived from infrared reverberation mapping studies (Kishimoto et al., 2007; Koshida et al., 2014). Following Figure 2 in Savić et al. (2018), the outer radius for the scattering region is chosen such that the BLR half-opening angle when viewed from the edge of the scattering region is 35o. This angle gives the angular extension of the BLR while keeping the observer at the outer edge of the scattering region. We choose a slightly larger value of the half-opening angle compared to Savić et al. (2018, where they assumed a BLR half-opening angle = 25o) which allows us to reproduce well the wings of the emission lines. This value was obtained after testing with a grid of values between 0o - 45o (with a step size of 5o; where 0o refers to no contribution from the BLR and 45o means that the outer edge of the BLR coincides with the outer edge of the scattering region). Values above 45o are discarded as they are nonphysical. Therefore, with the knowledge of the from the scaling relation of Netzer & Laor (1993), and this value of the BLR half-opening angle, we can estimate the outer radius for the scattering region ( = ). By combining information on inner and outer radii with electron number density, we estimate the optical depth of this region. A grid of electron number density solutions is explored to identify the optimal solution aligning with spectral data gathered from observations.
@@ 5 Discussions
We conducted spectro-polarimetric observations of the gamma-ray-emitting NLS1, 1H0323+342, renowned for its strong Fe II emission, with an R4570 of 1.14 and accretion at 37% of the Eddington limit. The continuum polarization is modest, measuring at only 0.12%, aligning with the observed low-polarization trend in other high-accreting NLS1 sources. Goodrich (1989) performed spectro-polarimetry of 18 NLS1s, majority of which show a very low level of polarization %. Robinson et al. (2011) utilized a two-component scattering model, determining that the inclination of a line-emitting accretion disk is not a principal parameter, suggesting the distinctive properties of these objects have a physical rather than geometrical origin. Indeed, recent studies on high-accreting NLS1 show higher metallicities in these objects (e.g., Panda et al., 2019, 2020; Śniegowska et al., 2021; Panda, 2021; Garnica et al., 2022; Floris et al., 2024).
@@ 5.4 Inferences about the inclination and geometry
The inter-percentile width mentioned in Table 4 can be used to calculate the inclination. The implies , and the disk thickness must be high enough to explain the low-inclination angle inferred from the STOKES modeling. Assuming the inclination angle degrees implies a = 0.8 or disk half-opening angle of 40 degrees (). A relatively higher inclination angle of 24 degrees will correspond to a H/R = 0.7 and, consequently, a disk half-opening angle of 35 degrees. Indeed, recent dynamical modeling of reverberation mapping data, e.g., Williams et al. (2018), supports the argument that BLR is a thick disk where many objects show a half-opening angle () larger than 40 degrees (here =90°corresponds to a sphere). The BLR can have flared disk-like geometry where increases with R as (see discussion in Collin et al., 2006). Such flaring ‘BOWL’-shaped BLR models can explain various observational signatures, including the mechanism responsible for producing Lorentzian profiles in Balmer lines in low-inclination systems dominated by turbulence (see Goad et al., 2012).
@@ Acknowledgement
We thank the referee for her/his valuable comments and suggestions, which helped to improve the quality of the manuscript. The authors thank Paul Smith for his valuable support in the observations and SPOL data reduction and Alessandro Capetti for his help in the estimation of inter-stellar polarization. JJ and SR acknowledge the partial support of SERB-DST, New Delhi, through SRG grant no. SRG/2021/001334. SP acknowledges the financial support of the Conselho Nacional de Desenvolvimento Científico e Tecnológico (CNPq) Fellowships 300936/2023-0 and 301628/2024-6. SP is supported by the international Gemini Observatory, a program of NSF NOIRLab, which is managed by the Association of Universities for Research in Astronomy (AURA) under a cooperative agreement with the U.S. National Science Foundation, on behalf of the Gemini partnership of Argentina, Brazil, Canada, Chile, the Republic of Korea, and the United States of America. NS acknowledges the financial support provided under the National Post-doctoral Fellowship (NPDF; Sanction Number: PDF/2022/001040) by the Science & Engineering Research Board (SERB), a statutory body of the Department of Science & Technology (DST), Government of India. Observations reported here were obtained at the MMT Observatory, a joint facility of the University of Arizona and the Smithsonian Institution.
Appendix A Supplementary figures